Reference to a Statement of Changes in Equity may assist in the calculation of the missing amount: Opening Owners Equity: 100,000 Closing Owners Equity: 145,000 Capital Introduced During Period: 7,000 Revenue:? Expenses: 62,000 Enter answer in XXXXXX format-do not use any symbols or punctuation. Answer: Check
Answer #1

Closing owners equity=Opening owners equity+Capital introduced during period+Net income

145,000=100,000+7000+Net income

Net income=145000-(100,000+7000)

=$38000

Net income=Revenue-Expenses

Revenue=(38000+62000)

=100,000